OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -4x - y < -6

OpenStudy (anonymous):

Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

Dashed line with shading on the side that includes the origin.

OpenStudy (anonymous):

Thank you! Would you mind maybe helping with a few more...?

OpenStudy (anonymous):

Sure:)

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 6 < 9 or x + 6 (is greater than or equal to) 15 x < 3 or x (is greater than or equal to) 9 A number line with an open circle on 3, a closed circle on 9, and shading in between. x 9 or x < 3 A number line with an open circle on 3, shading to the left, and a closed circle on 9, shading to the right.

OpenStudy (anonymous):

x + 6 > 15 x > 15 - 6 x > 9 x + 6 < 12 x < 12 - 6 x < 6 ANSWER : c and d are not equivalent....the reason being, is that there will be no closed circles because there is no equal signs.

OpenStudy (anonymous):

so the answer is a or b?

OpenStudy (anonymous):

also, Part 1: Solve the inequality -5x + 8 <= x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (unklerhaukus):

\[-5x + 8 \leq x - 4\] get all the \(x\) terms to one side, and all the other terms to the other side

OpenStudy (unklerhaukus):

what did you get for x?

OpenStudy (anonymous):

so it'd be like -5x + x <= 8 - 4? sorry, i'm not good with this stuff D;

OpenStudy (unklerhaukus):

hmm, not quite \[−5x+8≤x−4\] the steps i would suggest would be first to add \(5x\) to both sides of the inequality , and then add \(4\) to both sides of the inequality

OpenStudy (anonymous):

so... 12 <= 5x?

OpenStudy (unklerhaukus):

very close,

OpenStudy (anonymous):

12 <= -5x?

OpenStudy (unklerhaukus):

\[−5x+8≤x−4\\−5x+8+5x+4≤x−4+5x+4\\8+4≤x+5x\\12≤\]

OpenStudy (anonymous):

is it -5x?

OpenStudy (unklerhaukus):

x+5x=(1+5)x

OpenStudy (anonymous):

ohhhh, so 12 <= 6x? also, i think this is the last one, but Which of the following ordered pairs is not a solution to the inequality y < -x + 6? (-3, 0) (-1, 0) (4, 3) (3, -2)

OpenStudy (unklerhaukus):

yeah, 12 <= 6x now divide both sides by 6

OpenStudy (anonymous):

so it's basically 6 <= x?

OpenStudy (anonymous):

i also need help with that last question up there.

OpenStudy (unklerhaukus):

try that bit again

OpenStudy (anonymous):

what do you mean? so the answer isn't 6 <= x?

OpenStudy (anonymous):

i honestly just need the answer to the two questions because i'm about to run out of time D:

OpenStudy (unklerhaukus):

\[12 ≤6x\] dividing both sides by six \[12/6 ≤ 6x/6\]

OpenStudy (anonymous):

alright now what about this one Which of the following ordered pairs is not a solution to the inequality y < -x + 6? (-3, 0) (-1, 0) (4, 3) (3, -2)

OpenStudy (unklerhaukus):

what do you get when you simplified 12/6<=6x/6 ?

OpenStudy (anonymous):

6 <= x

OpenStudy (unklerhaukus):

well you got the right hand side right , but not the left hand side

OpenStudy (anonymous):

12 <= 6x

OpenStudy (anonymous):

but at this point i seriously just need the answer & i'll review the lesson later, i have about two minutes before my time is up

OpenStudy (unklerhaukus):

what is twelve divided by six?

OpenStudy (anonymous):

two sorry ^^

OpenStudy (unklerhaukus):

thats better!

OpenStudy (anonymous):

thanks xD but yeah, the last question, i really just need the answer to it ^^

OpenStudy (unklerhaukus):

to check a solution (x,y) for y < -x + 6 just substitute

OpenStudy (anonymous):

sub what, exactly?

OpenStudy (anonymous):

i've only got a minute left

OpenStudy (unklerhaukus):

try one of the four options

OpenStudy (unklerhaukus):

is 0 < -(-3) + 6 this a true statement?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answer c?

OpenStudy (unklerhaukus):

right

OpenStudy (anonymous):

thank you for all of your help, sorry for my confusion with that one question xD